A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1910–1919

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

10th [Ano18f]. 11th [Ano18f]. 1908 [Ste10]. 1917 [Ano18e]. 1918 [Ano19g].

9th [Ano18f].

Acre [Dyc14c]. Action [Ano16e, Can19]. Activities [Shi17b]. Additional [Shi17a]. Affecting [Rad19]. affinis [Car17]. Afford [Rad19]. After [Bow10b]. Against [Can19]. Alaska [Cob10]. Alewife [Fie14a]. Alkali [Eas18]. Alosa [Sny19]. amended [Ano19b]. American [Ano10a, Ano14a, Ano18e, Ano19g, Ano14b, Ano15a, Ano16a, Ano17a, Ano18a, Sea14]. americanus [Sul15]. Among [How14b]. Angling [Fea16]. Animal [Whi14]. Animals [Pri14b, She16]. Annual [Ano15e, Ano16g, Ano17f, Ano18f, Ano18e, Ano19g]. Aquarium [Mee14]. arenaria [Bel14]. Artificial [How17a]. Artificially [How17b]. Atlantic [Atk14]. August [Ano18e].

Back [Rad14b]. Basin [Buc14]. Bass [Bro16, Bur10, Car14, Cus17, Nes14, Ril17, Sny14, Sny16, Wor10]. Bay [Par16]. be [Ano18f]. Biological [Can18, Shi17b]. Black [Bro16, Cus17, Nes14]. Bluefish [Nic14]. Bodies [She14]. Box [Ano10f]. Breeding [Rei14a]. Brook [Mor18, Tho19]. Buffalo [Can18]. Buffalo-Fish [Can18]. Building [Mee14].

Called [Ale14]. Can [Bla16b]. Canned [Hov18]. Carolina [Bro16, Rad14a]. Carp [Bar10]. Carried [Emb18]. Cases [How14b]. Cat [Ale14]. Catalina [Hol15]. Catch [Buc14]. Catfish [Lea10, Shi17a, Shi17c]. Catfishes [RL16]. Census [Ste10]. Certain [Tho15a, Tho19]. Certificate [Ano14b, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a]. Changes [Gre16, Ano19b]. Changing [Nev19]. Channel [Shi17a, Shi17c]. Characters [She16]. Chars [Ken15]. Chesapeake [Par16]. City [Ano18f, Ano19g]. Clam [Bel14]. Co [Par16]. Co-Operation [Par16]. Coast [Cob14a]. Collecting [Can18]. Color [Pri16b]. Commercial [Cob18, How14b, Per16, Rad19]. Commission [Bul16]. Commissions [Sea14]. Committee [Ano10e, Ano10g]. Commonly [Ale14]. Concerning [Nic14, Rei14b]. Conditions [Bel14, Nev19, Pri14b]. Connecticut [Row19]. Conservation [Bla16b, Par16, Shi19, Sim17, Tul17]. Constitution [Ano10b, Ano14c, Ano15b, Ano16b, Ano17b, Ano18b, Ano19b]. Construction [Bro16]. Contributions [O’M18]. Control [Cob16a, Lip18]. Copper [Tit14]. Cornell [Emb18]. Crappie [Lea10]. Creosotes [Tho19]. Cultural [AHK19, Bro16, Shi17b, Ken15]. Culture [Buc10, Lau10, Mar16, Mee10a, Zal14]. Culturists [War15].

date [Ano19b]. Decline [Row19]. Declining [Cob16b]. Defects [Pri14a]. Department [Bul14]. Depths [She16]. Description [Sul15]. Desirable [Row14]. Destroying [Row16]. Destruction [Dow17, Tit14]. Details [Buc10]. Development [Rad18]. Diamond [Rad14b]. Diamond-Back [Rad14b]. Different [She16]. Discussion [Ano10c, Ano19c]. Disease [Bea10]. Diseases [Mac15]. Disposal [Mar16]. Distribution [Bro16, Lin14]. do [Bla16b]. Don’t [Ano18f]. during [Gre16].

Economic [Mar10]. Editorial [Ano15c, Ano16c, Ano16d, Ano17c, Ano18c, Ano19d, Osb18]. Effects [Tho15a, Tho19]. Efforts [Sny16]. Eggs [Can18, Pri14b, Tho15b]. Eighth [Ano19g]. Election [Ano14d, Ano15d, Ano17e, Ano18d, Ano19e]. Elimination [War18]. Embankments [Can19]. England [Dim16, Ken15]. Essentials [Lau10]. Executive [Ano10e]. Experiment [Ave18, Car14]. Experimental [She17, War15, Emb18]. Experiments [Emb18, Mee10a, Mor18, Zal14]. Extension [Pri19a]. Eyed [Tho15b].

Fair [Per16]. Fairport [Can18, Shi17b]. Family [Nev19]. Farmer [Ano16e].

Fasting [Mor16]. Federal [Cob16a, Lip18]. Feeding [Car14, Emb18, Mar14]. Fifth [Ano16g]. First [Dyc14a]. Fish [AH16, AHK19, Ale14, Ale16, Ano16e, Bar17, Bea10, Bla16b, Bow10b, Bul16, Can18, Dim14, Dyc14a, Dyc14b, Dyc14c, Eas18, Emb14, Han16, Hen19, Hol15, Hov18, Ken15, Lin14, Lin16, Lyd19, Mar16, Mer17, Mor18, Pri14a, Pri14b, Pri16a, Rei14b, Rob19b, Rob19c, Shi17b, Sim17, Tow14, Tul17, War15, Zal14, Sea14]. Fish-cultural [Ken15]. Fisherfolk [Smi14]. Fisheries [AH16, Ano10a, Ano14b, Ano14a, Ano15a, Ano16a, Ano17a, Ano18a, Ano18e, Ano19g, Bul14, Can18, Cob14a, Cob14b, Cob18, Dim16, Par16, Pri16a, Rad19, Sea14]. Fishery [Cob16b, Fie14a, Rad19, Ste10]. Fishes [Dow17, Hun19, Mac15, Rad18, Rei14a, Rei17, She14, She17, Tho15a, Tit14]. Fishing [Ave18, Lip18, Per16, Rei14a]. Fishway [Pri19b]. Flesh [Pri16b]. Flounder [Sul15]. fontinalis [Tho19]. Food [Dow17, Emb14, Fie16, Lyd19, Nev19, Rob19b, Row14, Row16, Shi17c]. Foreign [Ano10g]. Forest [Mer17]. Forget [Ano18f]. Form [Pri19b]. Forty [Ano15e, Ano16g, Ano17f, Ano18e, Ano19g]. Forty-Eighth [Ano19g]. Forty-Fifth [Ano16g]. Forty-Fourth [Ano15e]. Forty-Seventh [Ano18e]. Forty-Sixth [Ano17f]. Fourth [Ano15e]. Frederick [Mee10b]. Fresh [Cok16, How14b, How17b, Lyd19, Shi19]. Fresh-Water [Cok16, How17b, Shi19]. Freshwater [How14a]. Fry [Rob19a]. Future [Bar10].

Gambusia [Car17]. Game [Mer17]. Gas [She17]. General [Rip18]. Generation [How17b]. George [Mee10b]. Glochidia [How17a]. Gravel [Rob19a]. Grayling [Bic14]. Great [Row19, Dow17]. Greatest [Rad19]. Grounds [Gre16]. Growth [Bel14, Hun19, Shi17c].

Habits [Rei14a]. Halibut [Cob16b]. Hardening [Bow10b]. Hatchery [Rip18, Dyc14a, Dyc14b]. Hatching [Can18, Emb18, Rob19a, Wor10]. Health [Lin16]. held [Ano18f]. Herring [Dun18, How17a].

Ictalurus [Shi17a, Shi17c]. Illuminating [She17]. Importance [Tit17]. Improve [Hov18]. Improvement [Rei14a]. Inconsiderate [Hen19]. Incorporation [Ano14b, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a]. Increase [Bow10b]. Increasing [Buc18]. Indices [She14]. Indiscriminate [Hen19]. Industry [Fie14b, Row16, Row19]. Infection [How17a]. Influence [Mor16]. Inimical [Pri14b]. Inland [Tit17]. Institution [Smi14]. Instruction [Smi14]. Interest [Sea14, War15]. Interstate [AHK19, Cob16a]. Invasion [Tal10]. Investigations [Bel15]. Iowa [Can18]. Island [Row19]. italics [Ano19b].

Kansas [Dyc14a, Dyc14b]. King [Cob10]. Knowledge [Rei14a].

Laboratories [Rad19]. Lake [Eas18, AH16, Dun18]. Lakes [Tit14, Dow17].

Large [Bro16, Car14]. Larvae [Pri14b]. Laws [Tul17]. Leaping [War10]. Learning [Rei17]. Legislation [Shi19]. Library [Fea16]. Limit [Pri19a]. List [Ano10d, Ano15d, Ano16f, Ano17e, Ano18d, Ano19e]. Lists [Ano14d]. Lobsters [Mor16]. Louisiana [Tul17].

Marine [Fie16, She16, Whi14]. Markets [Rad18]. Maryland [Par16]. Massachusetts [Fie14a]. Meal [Emb14]. Meeting [Ano18f, Ano18e, Ano19g, Ano19f, Ano15e, Ano16g, Ano17f]. Members [Ano10d, Ano14d, Ano15d, Ano16f, Ano17e, Ano18d, Ano19e]. Membership [Ano14d, Ano15d, Ano17e, Ano18d, Ano19e]. Memorial [Mee10b]. Memoriam [Ano17d]. Metallic [Tho15a]. Method [Fos19]. Methods [Can18, Tul17]. Michigan [Bow10a, Han16]. Mile [Pri19a]. Minn [Ano18e]. Minnesota [Ave18]. Minnow [Car17]. Mississippi [Sim17, Lip18]. Mixing [Leo18]. Montana [Bic14]. Mouth [Bro16, Car14, Nes14, Cus17]. Mussel [Fie14b]. Mussels [Cok16, How14a, How14b, How17b, Lyd19, Shi19]. Mya [Bel14].

Narrowly [How14b]. National [Bla16a, Smi14]. Nature [RL16]. Necessity [Shi19]. Need [Smi14]. Neglected [Rad18, Tow14]. North [Rad14a]. Notes [AH16, Bic14, Bow10b, Bur10, Dyc14a, Fos19, Ken15, Nic15, Rad14a, Shi17a, Shi17c, Sny14, TL15, War10]. Nutrition [Mor18].

Obnoxious [Tit14]. Offered [O’M18]. Officers [Ano10e, Ano14e]. Oils [Tho19, Whi14]. One [Dyc14b, Hol15]. Oneida [AH16]. Operation [Par16]. Organizations [Dim14]. Outlook [Sny19]. Output [Buc18, Dyc14a]. Ova [Bow10b]. Overcome [Pri14a]. Oyster [Row19]. Oysters [Row14].

Pacific [Cob14a, Cob14b, Cob16b, Gre16, War10]. Paddle [Ale14, Ale16]. Paddle-Fish [Ale14, Ale16]. Palisades [AHK19]. Parasites [Lin14, Lin16]. Parasitism [How14b]. Parent [Rob19b]. Park [AHK19]. Pass [Pri14a]. Passes [Pri14a]. Past [Bar17]. Paul [Ano18e]. Peabody [Mee10b]. Pearl [How14a]. Pennsylvania [Bul14, Bul16]. Perch [Ano10c, Fos19]. Perfect [Pri14a]. Permanent [Tit17]. Philadelphia [Mee14]. Physiological [Gre16, She16]. Places [Ano19f]. Plant [War15]. Planting [Eas18, Hen19]. Play [Per16]. Plea [Bea10, Rei14b]. Poisoning [She17]. Policy [AHK19, Row16, Tit17]. Pollution [War18]. Polyodon [Ale14, Ale16]. Pond [Bro16, Can19, Dyc14c, Lau10, Tow14]. Ponds [Buc14, Tit14]. Porto [Nic15]. Possibilities [Dyc14c]. Possible [Sea14]. Potomac [Tal10]. Preliminary [AHK19, Bel15]. Present [Bar17]. Preservation [Cok16, Rei14b]. Presidents [Ano19f]. Private [Tow14]. Prizes [O’M18]. Problem [Bla16a]. Problems [Rad19]. Proceedings [Ano15e, Ano16g, Ano17f, Ano18e]. Proceeedings [Ano19g]. Producing [Row16]. Product [Hov18]. Production [Dow17]. Products [Cob18, Rad19]. Progress [Rad14b, Wor10]. Promise [Rad19]. Promote [Bla16b]. Propagation [Bro16, Cus17, Fos19, Lea10, Rad14b, Sim17]. Properties [Whi14]. Protecting [Tho10]. Protection [Can19, Hol15, Mer17, Tul17]. Pseudopleuronectes [Sul15]. Public [Lin16, Row16]. punctatus [Shi17a, Shi17c].

Quantitative [Gre16]. Quebec [Mer17]. Question [Ano10f].

Raceway [Rob19c]. Rainbow [Bow10a]. Rays [Rad14a]. Reared [How17b]. Rearing [How14a, Ril17, Shi17a, Shi17c, TL15]. Record [How14a]. Records [Rei14b]. Red [Pri16b]. Regulating [Bel14]. Relation [Mar14]. Relations [Per16, Ano10g]. Relief [Rad19]. Report [AHK19, Ano10g]. Reserve [Hol15]. Resource [Tow14]. Restricted [How14b]. Results [Emb18]. Rhode [Row19]. Rico [Nic15]. Ripen [Sny16]. River [How17a, Lip18, Sim17]. Routine [Rip18]. Run [Gre16].

Salmon [Atk14, Buc10, Cob10, Gre16, TL15, War10]. Salmons [Pri16b]. Salt [Tho15b]. Salts [Tho15a]. Salvelinus [Tho19]. Santa [Hol15]. sapidissima [Sny19]. Sea [Fie14b, Pri14b]. Seasonal [Lin14]. Second [How17b]. Senses [Rei17]. Separating [Tho15b]. September [Ano18f, Ano19g]. Service [Ano19f]. Seventh [Ano18e]. Sewage [Mar16]. Shad [Bla16a, Sny19]. Sharks [Rad14a]. Showing [Ano14d, Ano15d, Ano17e, Ano18d, Ano19e]. Sixth [Ano17f]. Size [Bow10b]. Small [Cus17, Nes14]. Small-Mouth [Nes14, Cus17]. Society [Ano18e, Ano19g, Ano14b, Ano15a, Ano16a, Ano17a, Ano18a, Cob14b, Sea14, Ano10a, Ano14a]. Some [Buc10, Emb18, Fea16, Gre16, How14b, Lau10, Mac15, Pri14a, Pri14b, Tul17, Rad14a, Whi14]. Sound [Row16]. Sources [Fie16]. South [Bro16]. spathula [Ale14, Ale16]. Spawning [Gre16]. Species [How14b]. Spines [RL16]. Spoonbill [Ale14]. Sport [Per16]. Sportsman [Mar10]. St [Ano18e]. Stages [Sul15]. State [Ave18, Bel15, Dyc14b, Shi19, War18]. States [Cob18]. Station [Bro16, Can18, Emb18, Shi17b]. Stocking [Bel15, Tit17]. Stray [Nic15]. Stream [War18]. Striped [Sny14, Sny16, Wor10]. Studies [Mor18]. Study [Bea10, She17, Tho19]. Sturgeon [Mee10a]. Suggested [Pri19a]. Suggestions [Sea14, She14, Pri14a]. Suitability [She14]. Sulphate [Tit14]. Superior [Dun18]. Supply [Rob19b]. Survey [Han16]. Systematic [Bea10, Bel15].

Tars [Tho19]. Taxonomic [Ken15]. Technical [Smi14]. Terms [Ano19f]. Terrapin [Rad14b]. Territorial [Pri19a]. Texas [Ril17]. their [Par16]. Them [Pri14a]. Three [Pri19a]. Thyroid [Mar14]. Top [Car17]. Trade [Dim14]. Transactions [Ano10a, Ano14a]. Trap [Rob19c]. Treasures [Fea16]. Trout [Bow10a, Emb14, Emb18, Leo18, Mar14, Mor18, Nev19, Rip18, Tho19, Tho10]. Trouts [Ken15, Pri16b]. Tumor [Mar14].

U.S. [Can18]. Ultimate [Fie16]. Ultization [Cok16]. Undersized [Tho10]. Unfertilized [Tho15b]. United [Cob18]. University [Emb18]. Unsolved [Rad19]. upon [Tho15a, Tho19]. Upper [Sim17]. Use [Tho15b, Tit14]. Utilization [Whi14].

Value [Mar10]. Virginia [Par16].

Walbaum [Sul15]. Waste [Bar17]. Wastes [She17]. Water [Bow10b, Cok16, How14b, How17b, Lyd19, She14, Shi19]. Waters [Bel15, Cob16a, Leo18, Pri19a, Tit17]. Wave [Can19]. Western [Leo18]. White [Fos19]. Winter [Sul15]. Work [Bul14, Bul16, Dyc14b].

Year [Ano14d, Ano15d, Ano17e, Ano18d, Ano19e, Dyc14a, Dyc14b, Hol15]. Yellow [Ano10c, Bur10]. York [Ano18f, Ano19g, War18]. Young [Car14, Nic14, Sul15].

Zealand [Pri16a].

References

Adams:1916:NOL

[AH16] Chas C. Adams and T. L. Hankinson. Notes on Oneida Lake fish and fisheries. Transactions of the American Fisheries Society, 45(3):154– 169, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1919:PRF

[AHK19] Chas C. Adams, T. L. Hankinson, and W. C. Kendall. A preliminary report on a fish cultural policy for the Palisades Interstate Park. Transactions of the American Fisheries Society, 48(4):193– 204, September 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alexander:1914:PFP

[Ale14] M. L. Alexander. The paddle-fish: (Polyodon spathula). (commonly called “spoonbill cat.”). Transactions of the American Fisheries Society, 44(1):73–78, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alexander:1916:MAP

[Ale16] M. L. Alexander. More about the paddle-fish (Polyodon spathula). Transactions of the American Fisheries Society, 45(1):34–39, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:TAF

[Ano10a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 39(1):9–58, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:C

[Ano10b] Anonymous. Constitution. Transactions of the American Fisheries Society, 39(1):221–223, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:DYP

[Ano10c] Anonymous. Discussion on yellow perch. Transactions of the American Fisheries Society, 39(1):192–195, January 1910. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:LM

[Ano10d] Anonymous. List of members, 1909–1910. Transactions of the American Fisheries Society, 39(1):197–219, January 1910. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:OEC

[Ano10e] Anonymous. Officers for 1909–1910 and Executive Committee. Transactions of the American Fisheries Society, 39(1):3, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:QB

[Ano10f] Anonymous. Question box. Transactions of the American Fisheries Society, 39(1):183–191, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1910:RCF

[Ano10g] Anonymous. Report of the Committee on Foreign Relations. Transactions of the American Fisheries Society, 39(1):174–182, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1914:TAF

[Ano14a] Anonymous. Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 43(1):7–35, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1914:CIA

[Ano14b] Anonymous. Certificate of incorporation of the American Fisheries Society. Transactions of the American Fisheries Society, 43(1):4, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1914:C

[Ano14c] Anonymous. Constitution. Transactions of the American Fisheries Society, 43(1):215–217, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1914:LMS

[Ano14d] Anonymous. Lists of members, 1913–1914: Showing year of election to membership. Transactions of the American Fisheries Society, 43(1):197–213, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1914:O

[Ano14e] Anonymous. Officers. Transactions of the American Fisheries Society, 43(1):2, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1915:CIA

[Ano15a] Anonymous. Certificate of incorporation of the American Fisheries Society. Transactions of the American Fisheries Society, 44(4):221, September 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1915:C

[Ano15b] Anonymous. Constitution. Transactions of the American Fisheries Society, 44(4):245–247, September 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1915:E

[Ano15c] Anonymous. Editorial. Transactions of the American Fisheries Society, 44(4):224–225, September 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1915:LMS

[Ano15d] Anonymous. List of members, 1914–1915: Showing year of election to membership. Transactions of the American Fisheries Society, 44 (4):226–244, September 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1915:PFF

[Ano15e] Anonymous. Proceedings of the Forty-Fourth Annual Meeting. Transactions of the American Fisheries Society, 44(3):179–214, June 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:CIA

[Ano16a] Anonymous. Certificate of incorporation of the American Fisheries Society. Transactions of the American Fisheries Society, 45(4):213, September 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:C

[Ano16b] Anonymous. Constitution. Transactions of the American Fisheries Society, 45(4):237–240, September 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:Ea

[Ano16c] Anonymous. Editorial. Transactions of the American Fisheries Society, 45(1):109–111, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:Eb

[Ano16d] Anonymous. Editorial. Transactions of the American Fisheries Society, 45(4):215–219, September 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:FFA

[Ano16e] Anonymous. The fish farmer in action. Transactions of the American Fisheries Society, 45(1):46–47, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:LM

[Ano16f] Anonymous. List of members, 1914–1915. Transactions of the American Fisheries Society, 45(4):220–236, September 1916. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1916:PFF

[Ano16g] Anonymous. Proceedings of the Forty-Fifth Annual Meeting. Transactions of the American Fisheries Society, 45(1):77–107, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:CIA

[Ano17a] Anonymous. Certificate of incorporation of the American Fisheries Society. Transactions of the American Fisheries Society, 46(4):202, September 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:C

[Ano17b] Anonymous. Constitution. Transactions of the American Fisheries Society, 46(4):223–226, September 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:E

[Ano17c] Anonymous. Editorial. Transactions of the American Fisheries Society, 46(2):125–126, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:M

[Ano17d] Anonymous. In memoriam. Transactions of the American Fisheries Society, 46(3):187–197, June 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:LMS

[Ano17e] Anonymous. List of members, 1916–1917: Showing year of election to membership. Transactions of the American Fisheries Society, 46 (4):205–221, September 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1917:PFS

[Ano17f] Anonymous. Proceedings of the Forty-Sixth Annual Meeting. Transactions of the American Fisheries Society, 46(3):171–186, June 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:CIA

[Ano18a] Anonymous. Certificate of incorporation of the American Fisheries Society. Transactions of the American Fisheries Society, 47(4):2, September 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:C

[Ano18b] Anonymous. Constitution. Transactions of the American Fisheries Society, 47(4):136–139, September 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:E

[Ano18c] Anonymous. Editorial. Transactions of the American Fisheries Society, 47(4):140–141, September 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:LMS

[Ano18d] Anonymous. List of members, 1917–1918: Showing year of election to membership. Transactions of the American Fisheries Society, 47 (4):119–135, September 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:PFS

[Ano18e] Anonymous. Proceedings of the Forty-Seventh Annual Meeting of the American Fisheries Society: St. Paul, Minn., August 29, 30 and 31, 1917. Transactions of the American Fisheries Society, 47(2):51–79, March 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1918:MDF

[Ano18f] Anonymous. The 1918 Meeting: Don’t Forget the Annual Meeting to be held in New York City, September 9th, 10th and 11th. Transactions of the American Fisheries Society, 47(3):114, June 1918. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:CI

[Ano19a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 48(4):205, September 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:CAD

[Ano19b] Anonymous. Constitution: (as amended to date, recent changes in italics). Transactions of the American Fisheries Society, 48(4):222– 225, September 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:D

[Ano19c] Anonymous. Discussion. Transactions of the American Fisheries Society, 48(2):136–138, March 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:E

[Ano19d] Anonymous. Editorial. Transactions of the American Fisheries Society, 49(1):42–43, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:LMS

[Ano19e] Anonymous. List of members, 1918–1919: Showing year of election to membership. Transactions of the American Fisheries Society, 48 (4):207–221, September 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:PTS

[Ano19f] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 48(4):206, September 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1919:PFE

[Ano19g] Anonymous. Proceeedings of the Forty-Eighth Annual Meeting of the American Fisheries Society: New York City September 9, 10 and 11, 1918. Transactions of the American Fisheries Society, 48(2):77–112, March 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1914:AS

[Atk14] Chas G. Atkins. The Atlantic salmon. Transactions of the American Fisheries Society, 43(1):173–177, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Avery:1918:MES

[Ave18] Carlos Avery. Minnesota’s experiment in state fishing. Transactions of the American Fisheries Society, 48(1):57–64, December 1918. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1910:FC

[Bar10] S. P. Bartlett. The future of the carp. Transactions of the American Fisheries Society, 39(1):151–154, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1917:FWP

[Bar17] S. P. Bartlett. Fish waste, past and present. Transactions of the American Fisheries Society, 47(1):22–27, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1910:PSS

[Bea10] Tarleton H. Bean. A plea for the systematic study of fish disease. Transactions of the American Fisheries Society, 39(1):65–78, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1914:CRG

[Bel14] David L. Belding. Conditions regulating the growth of the clam (Mya arenaria). Transactions of the American Fisheries Society, 43(1): 121–130, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1915:PIS

[Bel15] David L. Belding. Preliminary investigations for the systematic stocking of state waters. Transactions of the American Fisheries Society, 44(2):137–149, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bickford:1914:NMG

[Bic14] W. M. Bickford. Notes on the Montana grayling. Transactions of the American Fisheries Society, 43(1):153–155, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1916:SNP

[Bla16a] Charles Minor Blackford. The shad — a national problem. Transactions of the American Fisheries Society, 46(1):5–14, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1916:WWC

[Bla16b] Charles Minor Blackford. What we can do to promote fish conservation. Transactions of the American Fisheries Society, 45(1):13–18, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1910:RTM

[Bow10a] Seymour Bower. The rainbow trout in Michigan. Transactions of the American Fisheries Society, 39(1):130–142, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1910:NIS

[Bow10b] Ward T. Bower. Notes on the increase in size of fish ova after water hardening. Transactions of the American Fisheries Society, 39(1): 92–102, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1916:CPC

[Bro16] G. W. N. Brown. The construction of a pond cultural station, and the propagation and distribution of large mouth black bass in South Carolina. Transactions of the American Fisheries Society, 46(1):30– 34, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1910:SDS

[Buc10] W. O. Buck. Some details of salmon culture. Transactions of the American Fisheries Society, 39(1):120–123, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1914:CBP

[Buc14] W. O. Buck. Catch basin ponds. Transactions of the American Fisheries Society, 44(1):67–70, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1918:IO

[Buc18] W. O. Buck. Increasing the output. Transactions of the American Fisheries Society, 47(3):96–100, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buller:1914:WPF

[Bul14] N. R. Buller. Work of the Pennsylvania Fisheries Department. Transactions of the American Fisheries Society, 43(1):111–119, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buller:1916:WPF

[Bul16] N. R. Buller. The work of the Pennsylvania Fish Commission. Transactions of the American Fisheries Society, 45(3):145–153, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burnham:1910:NYB

[Bur10] Charles W. Burnham. Notes on the yellow bass. Transactions of the American Fisheries Society, 39(1):103–108, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Canfield:1918:MCH

[Can18] H. L. Canfield. Methods of collecting and hatching buffalo-fish eggs at the U.S. Fisheries Biological Station, Fairport, Iowa. Transactions of the American Fisheries Society, 47(3):85–88, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Canfield:1919:PPE

[Can19] H. L. Canfield. The protection of pond embankments against wave action. Transactions of the American Fisheries Society, 48(3):141– 145, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carter:1914:EFY

[Car14] E. N. Carter. An experiment in feeding young large mouth bass. Transactions of the American Fisheries Society, 44(1):71–72, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carter:1917:TMG

[Car17] E. N. Carter. The top minnow, (Gambusia affinis). Transactions of the American Fisheries Society, 46(2):101–105, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1910:KSA

[Cob10] John N. Cobb. The king salmon of Alaska. Transactions of the American Fisheries Society, 39(1):124–129, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1914:FPC

[Cob14a] John N. Cobb. The fisheries of the Pacific Coast. Transactions of the American Fisheries Society, 44(1):61–64, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1914:PFS

[Cob14b] John N. Cobb. The Pacific Fisheries Society. Transactions of the American Fisheries Society, 44(1):65–66, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1916:FCI

[Cob16a] Eben W. Cobb. Federal control of interstate waters. Transactions of the American Fisheries Society, 46(1):26–29, December 1916. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1916:PHF

[Cob16b] John N. Cobb. Pacific halibut fishery declining. Transactions of the American Fisheries Society, 45(3):129–136, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1918:PCF

[Cob18] John N. Cobb. Products of the commercial fisheries of the United States. Transactions of the American Fisheries Society, 48(1):71–74, December 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coker:1916:UPF

[Cok16] Robert E. Coker. The ultization and preservation of fresh-water mussels. Transactions of the American Fisheries Society, 46(1):39–49, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushman:1917:PSM

[Cus17] O. P. Cushman. Propagation of small-mouth black bass. Transactions of the American Fisheries Society, 46(2):113–116, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dimick:1914:FTO

[Dim14] Frederick F. Dimick. The fish trade organizations. Transactions of the American Fisheries Society, 43(1):157–159, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dimick:1916:NEF

[Dim16] Frederick F. Dimick. The New England fisheries, 1915. Transactions of the American Fisheries Society, 45(3):137–144, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downing:1917:PDF

[Dow17] S. W. Downing. Production and destruction of the food fishes of the Great Lakes. Transactions of the American Fisheries Society, 47(1): 28–35, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunn:1918:LSH

[Dun18] A. C. Dunn. The Lake Superior herring. Transactions of the American Fisheries Society, 47(3):92–95, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dyche:1914:NNK

[Dyc14a] L. L. Dyche. Notes on the new Kansas Fish Hatchery and the first year’s output. Transactions of the American Fisheries Society, 44(1): 5–12, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dyche:1914:OYW

[Dyc14b] L. L. Dyche. One year’s work at the Kansas State Fish Hatchery. Transactions of the American Fisheries Society, 43(1):77–85, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dyche:1914:PAF

[Dyc14c] L. L. Dyche. Possibilities of an acre fish pond. Transactions of the American Fisheries Society, 43(1):67–75, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eastgate:1918:PFA

[Eas18] Alfred Eastgate. Planting fish in an alkali lake. Transactions of the American Fisheries Society, 47(3):89–91, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1914:FMF

[Emb14] G. C. Embody. Fish meal as a food for trout. Transactions of the American Fisheries Society, 44(1):57–60, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1918:RST

[Emb18] G. C. Embody. Results of some trout feeding experiments carried on in the Experimental Hatching Station of Cornell University. Transactions of the American Fisheries Society, 48(1):26–33, December 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fearing:1916:ALS

[Fea16] Daniel B. Fearing. An angling library and some of its treasures. Transactions of the American Fisheries Society, 45(1):53–75, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1914:AFM

[Fie14a] G. W. Field. Alewife fishery of Massachusetts. Transactions of the American Fisheries Society, 43(1):143–151, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1914:SMI

[Fie14b] Irving A. Field. The sea mussel industry. Transactions of the American Fisheries Society, 43(1):131–142, January 1914. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1916:USM

[Fie16] Irving A. Field. The ultimate sources of marine food. Transactions of the American Fisheries Society, 45(3):177–190, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1919:WPN

[Fos19] Fred J. Foster. White perch notes and method of propagation. Transactions of the American Fisheries Society, 48(3):160–165, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:1916:SQP

[Gre16] C. W. Greene. On some quantitative physiological changes in the Pacific salmon during the run to the spawning grounds. Transactions of the American Fisheries Society, 45(1):5–12, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hankinson:1916:MFS

[Han16] T. L. Hankinson. A Michigan fish survey. Transactions of the American Fisheries Society, 46(1):35–38, December 1916. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1919:IIP

[Hen19] James Alexander Henshall. Indiscriminate and inconsiderate planting of fish. Transactions of the American Fisheries Society, 48(3):166– 169, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holder:1915:OYP

[Hol15] Charles F. Holder. One year of protection at the Santa Catalina Fish Reserve. Transactions of the American Fisheries Society, 44(2): 109–119, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hovden:1918:ICF

[Hov18] K. Hovden. Improve the canned fish product. Transactions of the American Fisheries Society, 47(3):103–104, June 1918. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howard:1914:NRR

[How14a] A. D. Howard. A new record in rearing freshwater pearl mussels. Transactions of the American Fisheries Society, 44(1):45–47, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howard:1914:SCN

[How14b] A. D. Howard. Some cases of narrowly restricted parasitism among commercial species of fresh water mussels. Transactions of the American Fisheries Society, 44(1):41–44, December 1914. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howard:1917:AIG

[How17a] A. D. Howard. An artificial infection with glochidia on the river herring. Transactions of the American Fisheries Society, 46(2):93– 100, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howard:1917:SGA

[How17b] A. D. Howard. A second generation of artificially reared fresh-water mussels. Transactions of the American Fisheries Society, 46(2):89–92, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1919:GF

[Hun19] A. G. Huntsman. The growth of fishes. Transactions of the American Fisheries Society, 49(1):19–23, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1915:TFC

[Ken15] William Converse Kendall. Taxonomic and fish-cultural notes on the chars or trouts of New England. Transactions of the American Fisheries Society, 44(2):97–108, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lauman:1910:SEP

[Lau10] Felix A. Lauman. Some essentials in pond culture. Transactions of the American Fisheries Society, 39(1):149–150, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1910:PCC

[Lea10] John L. Leary. Propagation of crappie and catfish. Transactions of the American Fisheries Society, 39(1):143–148, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leopold:1918:MTW

[Leo18] Aldo Leopold. Mixing trout in western waters. Transactions of the American Fisheries Society, 47(3):101–102, June 1918. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linton:1914:SDF

[Lin14] Edwin Linton. On the seasonal distribution of fish parasites. Transactions of the American Fisheries Society, 44(1):48–56, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linton:1916:FPP

[Lin16] Edwin Linton. Fish parasites and the public health. Transactions of the American Fisheries Society, 45(1):19–28, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lipinsky:1918:FCF

[Lip18] M. N. Lipinsky. Federal control of fishing in the Mississippi River. Transactions of the American Fisheries Society, 47(3):104–105, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1919:FWM

[Lyd19] D. Lydell. Fresh water mussels as a fish food. Transactions of the American Fisheries Society, 49(1):24–28, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacCallum:1915:SDF

[Mac15] G. A. MacCallum. On some diseases of fishes. Transactions of the American Fisheries Society, 44(2):154–158, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marlinton:1910:EVS

[Mar10] Andrew Price Marlinton. The economic value of the sportsman. Transactions of the American Fisheries Society, 39(1):109–119, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1914:FTR

[Mar14] M. C. Marsh. The feeding of trout in relation to thyroid tumor. Transactions of the American Fisheries Society, 44(1):13–19, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1916:SDF

[Mar16] M. C. Marsh. Sewage disposal by fish culture. Transactions of the American Fisheries Society, 45(3):170–176, June 1916. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1910:ESC

[Mee10a] W. E. Meehan. Experiments in sturgeon culture. Transactions of the American Fisheries Society, 39(1):85–91, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1910:MGF

[Mee10b] Wm. E. Meehan. Memorial: George Frederick Peabody. Transactions of the American Fisheries Society, 39(1):61–62, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1914:BAP

[Mee14] W. E. Meehan. Building an aquarium for Philadelphia. Transactions of the American Fisheries Society, 43(1):179–181, January 1914. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mercier:1917:FFG

[Mer17] Hon Honore Mercier. Fish, forest and game protection in Quebec. Transactions of the American Fisheries Society, 47(1):3–10, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgulis:1916:IFL

[Mor16] Sergius Morgulis. The influence of fasting on lobsters. Transactions of the American Fisheries Society, 45(3):198–201, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgulis:1918:SNF

[Mor18] Sergius Morgulis. Studies on the nutrition of fish: Experiments on brook trout. Transactions of the American Fisheries Society, 48(1): 34–56, December 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nesley:1914:SMB

[Nes14] Charles H. Nesley. Small-mouth black bass. Transactions of the American Fisheries Society, 43(1):165–167, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1919:CFC

[Nev19] James Nevin. Changing food conditions of the trout family. Transactions of the American Fisheries Society, 49(1):29–32, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nichols:1914:CYB

[Nic14] J. T. Nichols. Concerning young bluefish. Transactions of the American Fisheries Society, 43(1):169–172, January 1914. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nichols:1915:SNP

[Nic15] J. T. Nichols. Stray notes from Porto Rico. Transactions of the American Fisheries Society, 44(2):159–160, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:1918:POC

[O’M18] Henry O’Malley. Prizes offered for contributions. Transactions of the American Fisheries Society, 47(3):112–113, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osburn:1918:E

[Osb18] R. C. Osburn. Editorial. Transactions of the American Fisheries Society, 47(2):80–82, March 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1916:FCB

[Par16] Jno S. Parsons. The fisheries of the Chesapeake Bay and the cooperation of Maryland and Virginia for their conservation. Transactions of the American Fisheries Society, 46(1):15–25, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perce:1916:RCS

[Per16] H. Wheeler Perce. The relations of commercial and sport fishing — fair play. Transactions of the American Fisheries Society, 45(1): 29–33, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1914:PFP

[Pri14a] Edward E. Prince. A perfect fish pass: Some suggestions as to defects in fish passes and how to overcome them. Transactions of the American Fisheries Society, 43(1):47–56, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1914:SAC

[Pri14b] Edward E. Prince. Some animals and conditions inimical to fish eggs and larvae in the sea. Transactions of the American Fisheries Society, 43(1):57–66, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1916:FFN

[Pri16a] Edward E. Prince. The fish and fisheries of New Zealand. Transactions of the American Fisheries Society, 45(3):117–128, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1916:RCF

[Pri16b] Edward E. Prince. On the red color of the flesh in the salmons and trouts. Transactions of the American Fisheries Society, 46(1):50–61, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1919:TWS

[Pri19a] E. E. Prince. Territorial waters and a suggested extension of the Three Mile Limit. Transactions of the American Fisheries Society, 48(3):175–190, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1919:NFF

[Pri19b] Edward E. Prince. A new form of fishway. Transactions of the American Fisheries Society, 48(3):170–174, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1914:NSN

[Rad14a] Lewis Radcliffe. Notes on some North Carolina sharks and rays. Transactions of the American Fisheries Society, 44(1):37–40, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1914:PPD

[Rad14b] Lewis Radcliffe. Progress in the propagation of the diamond-back terrapin. Transactions of the American Fisheries Society, 44(1):33– 36, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1918:DMN

[Rad18] Lewis Radcliffe. The development of markets for neglected fishes. Transactions of the American Fisheries Society, 48(1):65–70, December 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1919:FPL

[Rad19] Lewis Radcliffe. Fishery products laboratories afford the greatest promise of relief of unsolved problems affecting commercial fisheries. Transactions of the American Fisheries Society, 49(1):3–18, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1914:IFT

[Rei14a] Jacob Reighard. Improvement of fishing through a knowledge of the breeding habits of fishes. Transactions of the American Fisheries Society, 43(1):97–103, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1914:PPR

[Rei14b] Jacob Reighard. A plea for the preservation of records concerning fish. Transactions of the American Fisheries Society, 43(1):105–110, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1917:SLF

[Rei17] Jacob Reighard. The senses and learning in fishes. Transactions of the American Fisheries Society, 46(3):133–170, June 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:1917:BRT

[Ril17] Mark Riley. Bass rearing in Texas. Transactions of the American Fisheries Society, 46(2):107–112, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ripple:1918:GRT

[Rip18] R. L. Ripple. The general routine of a trout hatchery. Transactions of the American Fisheries Society, 47(3):106–111, June 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1916:NSC

[RL16] H. D. Reed and T. J. Lloyd. The nature of the spines in catfishes. Transactions of the American Fisheries Society, 45(3):202–206, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1919:HFG

[Rob19a] A. Robertson. Hatching fry in gravel. Transactions of the American Fisheries Society, 48(3):146–156, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1919:PFF

[Rob19b] A. Robertson. The parent fish as a food supply. Transactions of the American Fisheries Society, 48(3):158–159, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1919:RFT

[Rob19c] A. Robertson. The raceway as a fish trap. Transactions of the American Fisheries Society, 48(3):157–158, June 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rowe:1914:ODF

[Row14] Henry C. Rowe. Oysters, a desirable food. Transactions of the American Fisheries Society, 43(1):183–185, January 1914. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rowe:1916:DFP

[Row16] Henry C. Rowe. Destroying a food producing industry: Is it sound public policy? Transactions of the American Fisheries Society, 46(1): 62–68, December 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rowe:1919:GDO

[Row19] Henry C. Rowe. The great decline of the oyster industry in Connecticut and Rhode Island. Transactions of the American Fisheries Society, 49(1):33–37, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seal:1914:SPI

[Sea14] Wm. P. Seal. Suggestions of possible interest to the American Fisheries Society and to Fish Commissions. Transactions of the American Fisheries Society, 43(1):187–194, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelford:1914:SIS

[She14] V. E. Shelford. Suggestions as to indices of the suitability of bodies of water for fishes. Transactions of the American Fisheries Society, 44(1):27–32, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelford:1916:PCM

[She16] V. E. Shelford. Physiological characters of marine animals from different depths. Transactions of the American Fisheries Society, 45 (3):191–197, June 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelford:1917:ESP

[She17] V. E. Shelford. An experimental study of the poisoning of fishes by illuminating gas wastes. Transactions of the American Fisheries Society, 46(2):73–75, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shira:1917:ANR

[Shi17a] Austin F. Shira. Additional notes on rearing the channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 47(1):45–47, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shira:1917:FCA

[Shi17b] Austin F. Shira. Fish cultural activities of the Fairport Biological Station. Transactions of the American Fisheries Society, 47(1):39– 44, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shira:1917:NRG

[Shi17c] Austin F. Shira. Notes on the rearing, growth, and food of the channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 46(2):77–88, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shira:1919:NSL

[Shi19] A. F. Shira. The necessity of State legislation in the conservation of fresh-water mussels. Transactions of the American Fisheries Society, 49(1):38–41, December 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpson:1917:CPF

[Sim17] Earl Simpson. Conservation and propagation of fish in the Upper Mississippi River. Transactions of the American Fisheries Society, 47(1):36–38, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1914:NNI

[Smi14] Hugh M. Smith. The need for a national institution for the technical instruction of fisherfolk. Transactions of the American Fisheries Society, 43(1):41–45, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1914:NSB

[Sny14] J. P. Snyder. Notes on striped bass. Transactions of the American Fisheries Society, 43(1):93–96, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1916:ERS

[Sny16] J. P. Snyder. Efforts to ripen striped bass, 1915. Transactions of the American Fisheries Society, 45(1):40–45, March 1916. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1919:SOA

[Sny19] J. P. Snyder. The shad outlook: (Alosa sapidissima). Transactions of the American Fisheries Society, 48(2):113–120, March 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevenson:1910:FC

[Ste10] Charles H. Stevenson. The Fishery Census of 1908. Transactions of the American Fisheries Society, 39(1):79–84, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sullivan:1915:DYS

[Sul15] W. E. Sullivan. A description of the young stages of the winter flounder: (Pseudopleuronectes americanus Walbaum). Transactions of the American Fisheries Society, 44(2):125–136, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbott:1910:IP

[Tal10] Henry Talbott. The invasion of the Potomac. Transactions of the American Fisheries Society, 39(1):168–173, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomson:1910:PUT

[Tho10] G. H. Thomson. Protecting the undersized trout. Transactions of the American Fisheries Society, 39(1):160–161, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1915:ECM

[Tho15a] Adrian Thomas. Effects of certain metallic salts upon fishes. Transactions of the American Fisheries Society, 44(2):120–124, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomson:1915:USS

[Tho15b] G. H. Thomson. The use of salt in separating unfertilized from eyed eggs. Transactions of the American Fisheries Society, 44(2):161–163, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1919:SEC

[Tho19] Adrian Thomas. A study of the effects of certain oils, tars, and creosotes upon brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 48(2):121–135, March 1919. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1914:UCS

[Tit14] John W. Titcomb. The use of copper sulphate for the destruction of obnoxious fishes in ponds and lakes. Transactions of the American Fisheries Society, 44(1):20–26, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1917:IPP

[Tit17] John W. Titcomb. Importance of a permanent policy in stocking inland waters. Transactions of the American Fisheries Society, 47(1): 11–21, December 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Torrey:1915:NRS

[TL15] H. B. Torrey and D. E. Lancefield. Notes on the rearing of salmon. Transactions of the American Fisheries Society, 44(2):150– 153, March 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Townsend:1914:PFP

[Tow14] C. H. Townsend. The private fish pond — a neglected resource. Transactions of the American Fisheries Society, 43(1):87–92, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tulian:1917:SLM

[Tul17] E. A. Tulian. Some of the laws and methods for the protection and conservation of Louisiana’s fish. Transactions of the American Fisheries Society, 46(2):117–123, March 1917. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1910:NLP

[War10] Henry B. Ward. Notes on the leaping of the Pacific salmon. Transactions of the American Fisheries Society, 39(1):162–167, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1915:EPI

[War15] Henry B. Ward. An experimental plant of interest to fish culturists. Transactions of the American Fisheries Society, 44(3):169–178, June 1915. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1918:ESP

[War18] Henry B. Ward. The elimination of stream pollution in New York State. Transactions of the American Fisheries Society, 48(1):3–25, December 1918. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:1914:PUS

[Whi14] George F. White. The properties and utilization of some marine animal oils. Transactions of the American Fisheries Society, 44(1): 79–91, December 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worth:1910:PHS

[Wor10] S. G. Worth. Progress in hatching striped bass. Transactions of the American Fisheries Society, 39(1):155–159, January 1910. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zalsman:1914:EFC

[Zal14] Philip G. Zalsman. Experiments in fish culture. Transactions of the American Fisheries Society, 43(1):161–163, January 1914. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).